OpenStudy (anonymous):

What is the solution to the equation 2^(1 + 2x )= 3 ?

OpenStudy (anonymous):

2^(1 + 2x )= 3 take the log_2 of both sides: 1+2x = log_2 3 subtract 1 from both sides: 2x = (log_2 3) - 1 divide by 2 x = [(log_2 3) - 1]/2

OpenStudy (anonymous):

-1/2 isnt an option....

OpenStudy (anonymous):

you didn't specify the way you wanted the solution... and -1/2 is not what I came up with. it's x = [(log_2 3) - 1]/2 did you want a calculator answer?

OpenStudy (anonymous):

umm, yea let me see what that looks like

OpenStudy (anonymous):

i dont have a calculator :(

OpenStudy (anonymous):

computer calculator?

OpenStudy (anonymous):

i have so many windows up lol

OpenStudy (turingtest):

just one more won't hurt http://www.wolframalpha.com

OpenStudy (anonymous):

i just need an answer to see if i did this right

OpenStudy (turingtest):

what did you get? I gave you a calculator...

OpenStudy (anonymous):

my computer will crash. its old i got .185

OpenStudy (turingtest):

http://www.wolframalpha.com/input/?i=%5B%28log_2+3%29+-+1%5D%2F2 if that doesn't work use a calculator from the applications section. Those have been there since time immemorial.

OpenStudy (anonymous):

is that the right answer

OpenStudy (anonymous):

0.292... is the "right" answer. If you want the most accurate answer you take the thing with the log.

OpenStudy (anonymous):

thank you very much. simple is all i need

OpenStudy (anonymous):

:) you are welcome